# Normal coactions extend to the C*-envelope

### 13 Mar 2024 v2

@@ 4.3. Groupoid-embeddable categories
Let $\mathfrak{G}$ be a discrete groupoid with range and source maps $\mathfrak{r}$ and $\mathfrak{s}$, respectively. By [BKQ21, Theorem 6.10], the universal group $\mathcal{U}(\mathfrak{G})$ of $\mathfrak{G}$ can be described as follows. Let $[u]\coloneq\mathfrak{r}(\mathfrak{s}^{-1}(u))$ denote the orbit of a unit $u\in\mathfrak{G}^{0}$ and let $\mathcal{R}$ be a complete set of representatives for the set of orbits $\mathfrak{G}^{0}/\mathfrak{G}\coloneq\{[u]:u\in\mathfrak{G}^{0}\}$, so that $\mathfrak{G}=\bigsqcup_{u\in\mathcal{R}}\mathfrak{G}_{[u]}$, where $\mathfrak{G}_{[u]}\coloneq\{g\in\mathfrak{G}:\mathfrak{s}(g)\in[u]\}$. For $u,v\in\mathfrak{G}^{0}$, we put $\mathfrak{G}_{u}^{v}\coloneq\{g\in\mathfrak{G}:\mathfrak{s}(g)=u,\mathfrak{r}(g)=v\}$, and for each $u\in\mathcal{R}$, we let $\mathcal{X}_{u}\coloneq[u]\setminus\{u\}$. Then, [BKQ21, Theorem 6.10] provides us with a group isomorphism
@@
$C_{r}^{*}(I_{l}\ltimes\Omega)\cong C_{0}(\Omega)\rtimes^{r}\mathcal{U}(\mathfrak{G})\quad\text{ and }\quad C_{r}^{*}(I_{l}\ltimes\partial\Omega)\cong C_{0}(\partial\Omega)\rtimes^{r}\mathcal{U}(\mathfrak{G}).$
@@ 5. Right LCM monoids
Recall that a monoid $P$ is right LCM if it is left cancellative and for all $c,d\in P$, we have that $cP\cap dP$ is either empty or of the form $rP$ for some $r\in P$. Right LCM monoids have been studied intensely by many operator algebraists, see, e.g., [St15, ABLS19, BLRS20, LOS21, LS22, NS22, St22, KKLL22b]. In this section (Theorem 5.4), we identify the C*-envelope of all right LCM monoids. This includes the class of gcd-monoids introduced by Dehornoy in [Doh17I, Doh17I], which contains concrete examples of monoids that are cancellative and right LCM, but not group-embeddable, see [DW17, Proposition 4.3]. Hence, our result applies to a large class of monoids that need not be group-embeddable, and are therefore not covered by previous works (for instance [Seh22]).
@@
If $[cd^{-1},\chi]=[ab^{-1},\chi]$, then there exists $X\in\chi^{-1}(1)$ such that $X\cap dP\cap bP\neq\emptyset$ and $cd^{-1}$ and $ab^{-1}$ agree on $X\cap dP\cap bP$. Thus, there exists $x,y,z\in P$ such that $z=dx=by$ and $cd^{-1}(z)=ab^{-1}(z)$, i.e., $cx=ay$. Now we have
@@
There is an isomorphism of topological groupoids $I_{l}^{0}\ltimes\partial\Omega\to\mathcal{U}(P_{0})\ltimes\partial\Omega$ given by $[cd^{-1},\chi]\mapsto(\rho_{0}(c)\rho_{0}(d)^{-1},\chi)$.
@@
For $[cd^{-1},\chi]\in I_{l}^{0}\ltimes\partial\Omega$, note that $(\rho_{0}(c)\rho_{0}(d)^{-1},\chi)=(\kappa_{0}(cd^{-1}),\mathfrak{s}((\rho_{0}(c)\rho_{0}(d)^{-1},\chi)))$, where $\kappa_{0}$ is the cocycle from Lemma 5.1. This shows that the map is well-defined; a short calculation shows that this map is indeed an isomorphism of topological groupoids. ∎
@@
The following resolves Li’s Question B for all cancellative right LCM monoids, including many non group-embeddable examples.
@@
Throughout this proof, we identify $\mathcal{A}_{r}(P)$ with its image in $C_{\textup{env}}^{*}(\mathcal{A}_{r}(P))$ under the canonical completely isometric embedding. Since $P$ is right LCM, the groupoid $I_{l}\ltimes\Omega$ is Hausdorff by [Li23, Corollary 4.2]. By [St22, Theorem 4.1] (see also [BS24, Theorem 2.1]) the canonical map $\pi_{\textup{env}}$ is injective if and only if the restriction $\varphi\coloneq\pi_{\textup{env}}|_{C^{*}(\partial\mathcal{A}_{r,0}(P))}$ is injective. Put
@@
[…]

### 17 Mar 2026 v3

@@ 4.3. Groupoid-embeddable categories
Let $\mathfrak{G}$ be a discrete groupoid with range and source maps $\mathfrak{r}$ and $\mathfrak{s}$, respectively. By [BKQ21, Theorem 6.10], the universal group $\mathcal{U}(\mathfrak{G})$ of $\mathfrak{G}$ can be described as follows. Let $u]\coloneq\mathfrak{r}(\mathfrak{s}^{-1}(u))$ denote the orbit of a unit $u\in\mathfrak{G}^{0}$ and let $\mathcal{R}$ be a complete set of representatives for the set of orbits $\mathfrak{G}^{0}/\mathfrak{G}\coloneq\{[u]:u\in\mathfrak{G}^{0}\}$, so that $\mathfrak{G}=\bigsqcup_{u\in\mathcal{R}}\mathfrak{G}_{[u]}$, where $\mathfrak{G}_{[u]}\coloneq\{g\in\mathfrak{G}:\mathfrak{s}(g)\in[u]\}$. For $u,v\in\mathfrak{G}^{0}$, we put $\mathfrak{G}_{u}^{v}\coloneq\{g\in\mathfrak{G}:\mathfrak{s}(g)=u,\mathfrak{r}(g)=v\}$, and for each $u\in\mathcal{R}$, we let $\mathcal{X}_{u}\coloneq[u]\setminus\{u\}$. Then, [BKQ21, Theorem 6.10] provides us with a group isomorphism
@@
[…]
@@ 5. Right LCM monoids
Recall that a monoid $P$ is right LCM if it is left cancellative and for all $c,d\in P$, we have that $cP\cap dP$ is either empty or of the form $rP$ for some $r\in P$. Right LCM monoids have been studied intensely by many operator algebraists, see, e.g., [BRRW14, St15, ABLS19, BLRS20, LOS21, LS22, NS22, St22, KKLL22b]. In this section (Theorem 5.4), we identify the C*-envelope of all right LCM monoids. This includes examples which are not group embeddable, including the class of Malcev type monoids from [EH25], as well as a class of gcd-monoids introduced by Dehornoy in [Doh17I, Doh17I] which contains examples which are not group-embeddable (see [DW17, Proposition 4.3]). Hence, our result applies to a large class of monoids that need not be group-embeddable, and are therefore not covered by previous works (for instance [Seh22]).
@@
If $cd^{-1},\chi]=[ab^{-1},\chi]$, then there exists $X\in\chi^{-1}(1)$ such that $X\cap dP\cap bP\neq\emptyset$ and $cd^{-1}$ and $ab^{-1}$ agree on $X\cap dP\cap bP$. Thus, there exists $x,y,z\in P$ such that $z=dx=by$ and $cd^{-1}(z)=ab^{-1}(z)$, i.e., $cx=ay$. Now we have
@@
There is an isomorphism of topological groupoids $I_{l}^{0}\ltimes\partial\Omega\to\mathcal{U}(P_{0})\ltimes\partial\Omega$ given by $cd^{-1},\chi]\mapsto(\rho_{0}(c)\rho_{0}(d)^{-1},\chi)$.
@@
For $cd^{-1},\chi]\in I_{l}^{0}\ltimes\partial\Omega$, note that $(\rho_{0}(c)\rho_{0}(d)^{-1},\chi)=(\kappa_{0}(cd^{-1}),\mathfrak{s}((\rho_{0}(c)\rho_{0}(d)^{-1},\chi)))$, where $\kappa_{0}$ is the cocycle from Lemma 5.1. This shows that the map is well-defined; a short calculation shows that this map is indeed an isomorphism of topological groupoids. ∎
@@
The following resolves Li’s Question B for all cancellative right LCM monoids, including many aforementioned non group-embeddable examples.
@@
Throughout this proof, we identify $\mathcal{A}_{r}(P)$ with its image in $C_{\textup{env}}^{*}(\mathcal{A}_{r}(P))$ under the canonical completely isometric embedding. Since $P$ is right LCM, the groupoid $I_{l}\ltimes\Omega$ is Hausdorff by [Nor14, Corollary 3.24] and [Ex08, Corollary 10.9].By [St22, Theorem 4.1] (see also [BS24, Theorem 2.1]) the canonical map $\pi_{\textup{env}}$ is injective if and only if the restriction $\varphi\coloneq\pi_{\textup{env}}|_{C^{*}(\partial\mathcal{A}_{r,0}(P))}$ is injective. Put
@@
$C^{*}(\partial\mathcal{A}_{r,0}(P))$$C^{*}(\mathcal{A}_{r,0}(P))$$C_{\textup{env}}^{*}(\mathcal{A}_{r,0}(P))$$C^{*}(\partial\mathcal{A}_{r,0}(P))^{\delta_{0}}_{e}$$C_{\textup{env}}^{*}(\mathcal{A}_{r,0}(P))^{\varepsilon_{\textup{env}}}_{e},$$\scriptstyle{E^{\delta_{0}}}$$\scriptstyle{\varphi}$$\scriptstyle{\psi}$$\scriptstyle{E^{\varepsilon_{\textup{env}}}}$$\scriptstyle{\psi\circ\varphi|_{C^{*}(\partial\mathcal{A}_{r,0}(P))^{\delta_{0}}_{e}}}$